OpenStudy (anonymous):

can you please show me how to solve? -(5/3)x+6=(3/2)x-1

OpenStudy (saifoo.khan):

\[-\frac 53x + 6 = \frac 32x - 1\] \[-\frac 53 x - \frac 32x = -1 - 6\] \[-\frac{19}{6}x= -7\] \[-19x = -42\] \[19x = 42\] \[x = \frac{42}{19}\]